Consider a long straight wire that carries a current of I1= 375 mA upward.

(a) What is the magnitude and direction of the magnetic field generated by the wire a distance of 1.00 cm to the left of the wire?

(b) Now consider that another long straight wire is placed in parallel with the first one, at the location considered in a, 1.00 cm to the left of the first wire. This new wire has a current of I2= 236.0 mA downward. What is the magnitude and direction of force per length of wire from the first wire on the second?

(c) Generalize your findings from b to make a statement about in which cases wires carrying current are attracted to, or repelled from, one another
Answer #1

Q B=Mo I 375 ( MA d and 375 mA = 375 x103 d=1 cm=10²m B=(4*X107) (375 X163) 2A(103) d=1 cm. B=150 xloo Tesla direction : Intoہے اس al cha alla Mo I, I2 2nd d= sepsation b/w wie = (4* X16+) (375 xios) (236 X163) 27 177,000 X10 f=1147 X10 6 N/m direct